Massachusetts Institute of Technology Handout 6 6.042J/18.062J: Mathematics for Computer Science September 17, 1998 Professors Nancy Lynch and Srinivas Devadas

Tutorial Assignments

Students in Section 1 (R01: 10-12 room 8-119) Instructor: Nancy Lynch/Srini Devadas

Michael Harder Ralph Harik Matt Hausmann Mark Huang Cynthia Johnson Matt Lee Vivian Ma Matt Reynolds Suzette Vandiver Michael Won Gina Yip

Students in Section 2 (R02: 10-12 room 2-142) Instructor: Rodrigo Leroux

Henry Atkins Christopher Avrich Luke Douglas Ashima Dua Nathan Ickes Michael Newton Matt Norwood Archit Shah Richard Su

Students in Section 3 (R03: 11-1 room 26-210) Instructor: Brian Dean

Ambreen Amjad Emily Chang Joseph Cohen Andrew Copeland Michael Ferraro Joshua Furman Megan Henry Ajay Kulkarni Jorge Lopez John Paul Strachan Shane Swenson Andre Van Horn

Students in Section 4 (R04: 1-3 room 26-168) Instructor: Brad Banks

Pearlin Cheung Neil Chungfat Nick Feamster Jeff Klann Ashwin Krishnamurthy Chak Kwok Jimmy Lin Yi-Fung Lin Eric Mitchell John Morris Alyssa Robinson Anne Thompson

Students in Section 5 (R05: 1-3 room 36-839) Instructor: Rodrigo Leroux

Zachary Bodner Laura Currea Russell Miller Omprakash Gnawali Geoffrey Goodell Dalie Jimenez Albert Leija Mandy Mobley Harry Portlock II Francis St.Louis Zachary Warren Nina Ye Brent Yen Misha Zitser

Students in Section 6 (R06: 2-4 room 26-310 Instructor: Brian Dean

Ken Conley Blair Dunn Julie Hsieh Derek Lam Roxanne Lau Usman Mobin Terry Poon Sunil Rao Steve Song Clint Sieunarine Joseph Tsiang Jon Xu

Students in Section 7 (R07: 2-4 room 24-402) Instructor: Sarfraz Khurshid

Ya-Bing Chu Shirley Hao Ramon Romero Y Vigil Daniel Good Ed Solovey Ravindra K. Sarma Kenneth Grant Follansbee Daniel R. Gomez Xinyu Hugo Liu Alexandar Yip Kalpak Kothari Bhuvana Kulkarni Ali Ibrahim

Students in Section 8 (R08: 10-12 room 36-372) Instructor: Sarfraz Khurshid

Paul Thordarson David Robison Marcin Strojwas Peter Gamache Danielle Hinton Kennet Belenky Wesley Chan Darren Collins Mu’taz Qubbaj

Students in Section 9 (R09: 2-4 room 24-407) Instructor: Adam Gould

Jordan Adler Chad Brown Yamel Cuevas Maik Flanagin Walter Holland Peter Litwack Emily Liu Sneha Thakkar Edward Tolson

Students in Section 10 (R10: 2-4 room 36-144) Instructor: John Dunagan

Archit P Shah Michael Tsai Fred Fagerstrom Manu Sridharan Rami Musa Amay N Champaneria Matthew S. Munsey Shantanu K Sinha Jamie Colin Rasmussen Arun A Tharuvai Tina K Tyan Aaron D Adler Pauline K. Hsu Andy Nakrin

Students in Section 11 (R11: 9-11 room 26-168) Instructor: Alex Duran

Xing Gu Bennett Landman Wanchung Cheung Mark Ethier Jessica Yeh Misha Koshelev Ian Finn Michael Mirvis Tarjei Mikkelsen Frank Cortez Julie Hwang Jodi Beggs

Students in Section 12 (R12: (3-5 room 26-168 Instructor: Brad Banks

William Abernathy Tarik Alatovic Jeff Billing James Chen Mansim Cheng Ben Fu Dharmesh Mehta Ripal Nathuji Jim O’Donnell Chris Porter Paul Rafiuly Kin-Joe Sham Andreas Sundquist

Students in Section 13 (R13: 11-1 room 2-139 Instructor: Alex Duran

Zhao Qing Dong Alan Asbeck Orton Huang Nori Yoshida Jennifer Maurer Benjamin Morse Pietro Russo Philip Rolfe Christopher Cheng Shastri Sandy Minpont Chien Crystal Harris

Students in Section 16 (R16: 11-1 room 2-147 Instructor: Adam Gould

Seth Bisen-Hersh Cynthia M Chow Todd Hiers Tony Jimenez Jocelyn Lin Matt Malchano Anthony Pelosi Simone Pugh Natalie Skorodinsky Peter Ree Bill Thies Allison Waingold